# Decaying turbulence beneath surface waves

### v1

@@ 4 Discussion
In this paper we point out the similarity between rotating turbulence on the beta-plane and turbulence beneath surface waves. In particular, turbulence beneath surface waves exhibits the formation of coherent structures, the development of zonal jets, and the suppression of kinetic energy dissipation. These features are consistent with known properties of turbulence beneath surface waves, but the connection with rotating turbulence is obscured in the vast majority of studies that also involve surface wind stress, ambient Lagrangian-mean shear, and surface forcing.
@@
This phenomenological model hints at a new way to understand how surface waves enhance turbulent mixing. In the paradigm proposed by McWilliams et al. (1997), the effect of surface waves on turbulence is associated with the “Stokes contribution” to shear production in the turbulent kinetic energy budget. However, our results show that this interpretation must be incomplete, because surface waves also control the evolution of initially shear-free flows. We suggest that the impact of surface waves may be instead linked to their tendency to catalyze, without exchanging energy, an increase in the correlation times and length scales of turbulent motions. A benefit to this interpretation is that turbulent shear production can be interpreted in the standard way: as a transfer of kinetic energy from the horizontally-averaged, Lagrangian-mean velocity, whose energy is otherwise conserved. In other words, our descriptive analysis leads to an alternative paradigm for wave-modified turbulence wherein shear production is “unchanged” relative to wave-free turbulence (as in rotating turbulence). Instead of modifying shear production directly, then, waves impact both mixing and kinetic energy by inducing an inverse cascade, increasing the turbulent mixing length, and suppressing kinetic energy dissipation.
@@ 4.1 The “pseudovorticity” number
The analogy with rotating turbulence leads to a new non-dimensional number for characterizing surface wave effects on turbulence. In rotating turbulence, the Rossby number measures the relative magnitude of the relative vorticity ×$\bm{\nabla}\times\bm{u}$ and the background vorticity, $f\hskip 1.00006pt\bm{\hat{z}}$, such that
@@
where $W$ is a vertical velocity scale, $\Omega$ is the magnitude of the Stokes drift, and $H$ is a vertical length scale. The pseudovorticity number $\mathrm{Ps}$ measures the average role of surface waves with Stokes shear magnitude $\Omega$ on turbulent motions with vertical velocity scale $W$ and turbulent scale $H$. In the two-equation model in (10)–(11), $\mathrm{Ps}\sim\epsilon/k\Omega$. In our decaying scenarios, $\mathrm{Ps}$ eventually vanishes at $t\to\infty$ and dissipation is completely suppressed.
@@ 4.2 Connection with the Langmuir number
where $u_{\star}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sqrt{\tau}$ is the friction velocity, which is the square root of the kinematic wind stress $\tau$. At face value $\mathrm{La}$ does not apply to decaying cases, but we can amend this by interpreting $u_{\star}$ more generally as a turbulent velocity scale.
@@
[Acknowledgements]Without implying their endorsement, we would like to acknowledge fruitful discussions with Bruno Deremble, Petros Ioannou, and Bill Young.
@@
[Funding] G.L.W. was supported by the National Science Foundation, grant OCE-2342715. N.C.C. was supported by the Australian Research Council under DECRA Fellowship DE210100749, by the Center of Excellence for the Weather of the 21st Century CE230100012, and the Discovery Project DP240101274.
@@
[Data availability statement]Scripts to reproduce the simulations and analyses in this of this study are openly available at GitHub repository github.com/glwagner/WaveAveragedDecayingTurbulence.

### v2

@@ 4 Discussion
In this paper we point out the similarity between rotating turbulence on the beta-plane and turbulence beneath surface waves. In particular, turbulence beneath surface waves exhibits the formation of coherent structures perpendicular to the direction of wave propagation, the development of zonal jets, and the suppression of kinetic energy dissipation. These features are consistent with known properties of turbulence beneath surface waves, but the connection with rotating turbulence is obscured in studies that also involve surface wind stress, ambient Lagrangian-mean shear, and surface forcing. In particular, we hypothesize that some of the generic features of Langmuir turbulence — such as the formation of coherent structures, a suppression of kinetic energy dissipation, and an increase in mixing — are driven by similar dynamics as the phenomenon observed in our decaying situations. At the same time, however, the presence of Lagrangian-mean shear produces additional and distinctly different phenomenon, most clearly in the alignment of coherent vortices with the Lagrangian-mean shear, rather than perpendicular to the direction of wave propagation.
@@
This phenomenological model hints at a new way to understand how surface waves enhance turbulent mixing. In the paradigm proposed by McWilliams et al. (1997), the effect of surface waves on turbulence is associated with the “Stokes contribution” to shear production in the turbulent kinetic energy budget. However, our results show that this interpretation must be incomplete, because surface waves also control the evolution of initially shear-free flows.
We therefore propose an alternative theory which supposes that the impact of surface waves may be instead linked to their tendency to catalyze, without exchanging energy, an increase in the correlation times and length scales of turbulent motions. A benefit to our alternative interpretation is that turbulent shear production can be interpreted in the standard way: as a transfer of kinetic energy from the horizontally-averaged, Lagrangian-mean velocity, whose energy is otherwise conserved. In other words, our descriptive analysis leads to an alternative paradigm for wave-modified turbulence wherein shear production is “unchanged” relative to wave-free turbulence (as in rotating turbulence). Instead of modifying shear production directly, then, waves impact both mixing and kinetic energy by inducing an inverse cascade, increasing the turbulent mixing length, and suppressing kinetic energy dissipation.
@@ 4.1 The “pseudovorticity” number
The analogy with rotating turbulence inspires the definition of a new non-dimensional number for characterizing surface wave effects on turbulence. In rotating turbulence, the Rossby number measures the relative magnitude of the relative vorticity ×$\bm{\nabla}\times\bm{u}$ and the background vorticity, $f\hskip 1.00006pt\bm{\hat{z}}$, such that
@@
where $W$ is a turbulence velocity scale, $\Omega$ is the magnitude of the Stokes drift shear, and $H$ is a turbulence length scale. In boundary layer turbulence, for example, $W$ and $H$ are most straightforwardly measured by the magnitude of the vertical velocity and the boundary layer depth.
The pseudovorticity number $\mathrm{Ps}$ measures the average role of surface waves with Stokes shear magnitude $\Omega$ on turbulent motions with turbulent velocity scale $W$ and turbulent length scale $H$. In the two-equation model in (10)–(11), $\mathrm{Ps}\sim\epsilon/k\Omega$. In our decaying scenarios, $\mathrm{Ps}$ eventually vanishes at $t\to\infty$ and dissipation is completely suppressed.
@@ 4.2 Connection with the Langmuir number
where $u_{\star}$ is the friction velocity, which is the square root of the kinematic wind stress. At face value, $\mathrm{La}$ does not apply to decaying cases, but we can amend this by interpreting $u_{\star}$ more generally as a turbulent velocity scale.
@@
[Acknowledgements]Without implying their endorsement, we would like to acknowledge fruitful discussions with Bruno Deremble, Petros Ioannou, and Bill Young. We would also like to thank the editor and the three anonymous reviewers for their constructive comments that greatly improved our paper. Part of this work was done during the Sean R. Haney Memorial Symposium 2025 at the Scripps Institution of Oceanography in La Jolla, California.
@@
[Funding] G.L.W. was supported by the National Science Foundation, grant OCE-2342715. N.C.C. was supported by the Australian Research Council under the DECRA Fellowship DE210100749, the Center of Excellence for the Weather of the 21st Century CE230100012, and the Discovery Project DP240101274.
@@
[Data availability statement]Scripts to reproduce the simulations and analyses in this of this study are openly available at the GitHub repository github.com/glwagner/WaveAveragedDecayingTurbulence.
Simulation output used to produced the figures in this paper will be uploaded in a Zenodo repository upon acceptance.